OpenStudy (anonymous):

Point A is the origin. Find the Coordinates of the two points that are 5 units and the x coordinate should be 3. (Hint: Make a sketch. The two points fall on the line x=3 and therefore have coordinates (3,y) where y is unknown. Use the distance formula suitably rearranged.) A(0,0) B(3,y) X1=0 X2=3 y1=0 y2=y Should I use the distance formula? d=√(x2-x1)^2+(y2-y1)^2

OpenStudy (amistre64):

you can, which is just the pythag thrm rearranged

OpenStudy (anonymous):

right but It doesn't make sense using distance on this, or does it

OpenStudy (amistre64):

it does ... distance is the length of the hypotenuse of a right triangle

OpenStudy (amistre64):

a^2 + b^2 = c^2 sqrt(a^2 + b^2) = c x^2 + y^2 = d^2 sqrt(x^2 + y^2) = d

OpenStudy (anonymous):

I have the D it's 5

OpenStudy (amistre64):



OpenStudy (mathstudent55):